Find the magnitude and direction of (a) the resultant (b) the equilibrium of two forces, 10N acting in the direction N30east and 15 N acting in the easterly direction, if both forces act at a point
1
Expert's answer
2020-12-08T10:52:42-0500

1)


"\vec{R}=(10\cos{30},10\sin{30})+(15,0)\\=(23.66,5)\ N\\R=\sqrt{23.66^2+5^2}=24.2\ N\\
\theta=\arctan{\frac{5}{23.66}}=11.9\degree\text{ N of E}"

2)


"U=R=24.2\ N\\
\alpha=11.9\degree\text{ S of W}"